OpenStudy (lovelyharmonics):

When 3.5 mol Al react with 12.5 mol HCl, what is the limiting reactant and how many moles of H2 can be formed? 2 Al + 6 HCl “yields”/ 2 AlCl3 + 3 H2

OpenStudy (lovelyharmonics):

i only have math problems left now .-. im no good at math

ganeshie8 (ganeshie8):

2Al require 6HCl that means, 1 Al require 3 HCl, right ?

OpenStudy (lovelyharmonics):

yus

ganeshie8 (ganeshie8):

you got 3.5 mol Al, so lets see if we have sufficient HCl for this : 1 Al require 3 HCl 3.5Al require 3.5*3 HCl

OpenStudy (lovelyharmonics):

so itd require 7.0 HCl

ganeshie8 (ganeshie8):

whats the value of 3.5*3 ?

OpenStudy (lovelyharmonics):

so al would be the limiting reactant since we dont have enough of it

ganeshie8 (ganeshie8):

whats the value of 3.5*3 ?

OpenStudy (lovelyharmonics):

i already said XD itd be 7.0 HCl

ganeshie8 (ganeshie8):

check again :o

ganeshie8 (ganeshie8):

3.5*3 = 3.5 + 3.5 + 3.5 = ?

OpenStudy (lovelyharmonics):

oh itd be 10.5

ganeshie8 (ganeshie8):

yes :) but still you're right about the limiting reactant : Al is the limiting reactant because it will get used up first leaving 12.5-10.5 = 2mol of HCl unused

OpenStudy (lovelyharmonics):

so then 3.5 mols of H2 owuld be formed

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

3.5mol of Al is getting reacted, lets find out how many mols of H2 will be foremed

ganeshie8 (ganeshie8):

given equation : `2 Al` + 6 HCl “yields”/ 2 AlCl3 + `3 H2`

ganeshie8 (ganeshie8):

Notice that, 2Al is producing 3H2 that means, 1Al will produce 3/2 H2, right ?

OpenStudy (lovelyharmonics):

yes...

ganeshie8 (ganeshie8):

that also means, 3.5 Al will produce 3.5*(3/2) H2, yes ?

OpenStudy (lovelyharmonics):

yes, thad mean that 5 1/4 would be produced

ganeshie8 (ganeshie8):

Correct !